Use the approximation that for each time step. You throw a metal block of mass 0.27 kg into the air, and it leaves your hand at time 0 at location <0, 2, 0> m with velocity <3.5, 4.0, 0> m/s. At this low velocity air resistance is negligible. Using the iterative method shown in Section 2.4 with a time step of 0.05 s, calculate step by step the position and velocity of the block at 0.05 s, 0.10 s, and 0.15 s. (Express your answers in vector form.) 0.05 s < , , > m 0.05 s < , , > m/s 0.10 s < , , > m 0.10 s < , , > m/s 0.15 s < , , > m 0.15 s < , , > m/s
Answer #1

An initial position (x0) in vector form which will be given as -

x0 = (2 m) \hat{}\hat{j}

An intial velocity (v0) in vector form which will be given as -

v0 = (3.5 m/s) \hat{i} + (4 m/s) \hat{j}

(i) The position of a metal block after time (t) which will be given by -

x = x0 + v0 t + (1/2) a t2

At t = 0.05 s, we get

x = (2 m) \hat{}\hat{j} + [(3.5 m/s) \hat{i} + (4 m/s) \hat{j}] (0.05 s) - [(0.5) (9.8 m/s2) \hat{j} (0.05 s)2]

x = (2 m) \hat{}\hat{j} + (0.175 m) \hat{i} + (0.2 m) \hat{j} - (0.01225 m) \hat{j}

x = (0.175 m) \hat{i} + (2.187 m) \hat{j}

The velocity of a metal block after time (t) which will be given by -

v = v0 + a t

At t = 0.05 s, we get

v = [(3.5 m/s) \hat{i} + (4 m/s) \hat{j}] - (9.8 m/s2) \hat{j} (0.05 s)

v = (3.5 m/s) \hat{i} + (4 m/s) \hat{j} - (0.49 m/s) \hat{j}

v = (3.5 m/s) \hat{i} + (3.51 m/s) \hat{j}

(ii) The position of a metal block after time (t) which will be given by -

x = x0 + v0 t + (1/2) a t2

At t = 0.10 s, we get

x = (2 m) \hat{}\hat{j} + [(3.5 m/s) \hat{i} + (4 m/s) \hat{j}] (0.10 s) - [(0.5) (9.8 m/s2) \hat{j} (0.10 s)2]

x = (2 m) \hat{}\hat{j} + (0.35 m) \hat{i} + (0.4 m) \hat{j} - (0.049 m) \hat{j}

x = (0.35 m) \hat{i} + (2.35 m) \hat{j}

The velocity of a metal block after time (t) which will be given by -

v = v0 + a t

At t = 0.10 s, we get

v = [(3.5 m/s) \hat{i} + (4 m/s) \hat{j}] - (9.8 m/s2) \hat{j} (0.10 s)

v = (3.5 m/s) \hat{i} + (4 m/s) \hat{j} - (0.98 m/s) \hat{j}

v = (3.5 m/s) \hat{i} + (3.02 m/s) \hat{j}

(iii) The position of a metal block after time (t) which will be given by -

x = x0 + v0 t + (1/2) a t2

At t = 0.15 s, we get

x = (2 m) \hat{}\hat{j} + [(3.5 m/s) \hat{i} + (4 m/s) \hat{j}] (0.15 s) - [(0.5) (9.8 m/s2) \hat{j} (0.15 s)2]

x = (2 m) \hat{}\hat{j} + (0.525 m) \hat{i} + (0.6 m) \hat{j} - (0.11025 m) \hat{j}

x = (0.525 m) \hat{i} + (2.489 m) \hat{j}

The velocity of a metal block after time (t) which will be given by -

v = v0 + a t

At t = 0.15 s, we get

v = [(3.5 m/s) \hat{i} + (4 m/s) \hat{j}] - (9.8 m/s2) \hat{j} (0.15 s)

v = (3.5 m/s) \hat{i} + (4 m/s) \hat{j} - (1.47 m/s) \hat{j}

v = (3.5 m/s) \hat{i} + (2.53 m/s) \hat{j}